# LancBiO: Dynamic Lanczos-aided Bilevel Optimization via Krylov Subspace

### v1

@@ 2.3 Relation to the existing algorithms
$v_{k}$ $=v_{k-1}-\eta\left(A_{k}v_{k-1}-b_{k}\right)$
$=(I-\eta A_{k})v_{k-1}+\eta b_{k},$
while the proposed SubBiO constructs a two-dimensional subspace, $\mathcal{S}_{k}=\operatorname{span}\left\{b_{{k}},(I-\eta A_{{k}})v_{k-1}\right\}$, defined in (8). It is worth noting that the updated $v_{k}$ in SOBA and FSLA belongs to the subspace $\mathcal{S}_{k}$. Furthermore, in the sense of solving the two-dimensional subproblem (6), SubBiO selects the optimal solution $v$ in the subspace.
In addition, if the subspace dimension $m$ is set to one, LancBiO is simplified to a scenario in which one conjugate gradient (CG) step with the warm start mechanism is performed in each outer iteration, which exactly recovers the algorithm AmIGO-CG (Arbel & Mairal, 2022) with one inner iteration to the update of $v$. Alternatively, if the step size $\lambda$ in Algorithm 2 is set to 00 within each $m$-steps (i.e., only inner iterations are invoked), LancBiO reduces to the algorithm AmIGO-CG (Arbel & Mairal, 2022) with $m$ inner iterations.
@@ 3 Theoretical Analysis
The Lipschitz properties of $f,g$ and the strong convexity of the lower-level problem revealed by the above assumptions are standard in bilevel optimization (Ghadimi & Wang, 2018; Chen et al., 2021; Ji et al., 2021; Khanduri et al., 2021; Arbel & Mairal, 2022; Chen et al., 2022; Dagréou et al., 2022; Li et al., 2022; Ji et al., 2022; Hong et al., 2023). These assumptions ensure the smoothness of $\varphi$ and $y^{*}$; see the following results (Ghadimi & Wang, 2018).
@@
Under the Assumptions 3.2 and 3.3, $y^{*}(x)$ is ${L_{gx}}/{\mu_{g}}$Lipschitz continuous, i.e., for any $x_{1},x_{2}\in\mathbb{R}^{d_{x}}$,
$\left\|{y^{*}(x_{1})-y^{*}(x_{2})}\right\|\leq\frac{L_{gx}}{\mu_{g}}\left\|{x_{1}-x_{2}}\right\|.$ (11)
@@
Under the Assumptions 3.1, 3.2 and 3.3, the hyper-gradient $\nabla\varphi(\cdot)$ is $L_{\varphi}$-Lipschitz continuous, i.e., for any $x_{1},x_{2}\in\mathbb{R}^{d_{x}}$,
where $L_{\varphi}>0$ is defined in Appendix E.
@@
3.6, commonly adopted in (Ghadimi & Wang, 2018; Ji et al., 2021; Liu et al., 2022; Kwon et al., 2023b), is helpful in ensuring the stable behavior of the dynamic Lanczos process; see Section 3.1.
@@ 3.1 Subspace Properties in Dynamic Lanczos Process
In view of the inherent instability of the Lanczos process (Paige, 1980; Meurant & Strakoš, 2006) and the evolution of the Hessian $\{A_{k}\}$ and the gradient $\{b_{k}\}$ in LancBiO, the analysis of the constructed subspaces is intricate. Based on the existing work (Paige, 1976, 1980; Greenbaum, 1997), this subsection sheds light on the characterization of subspaces and the effectiveness of the subproblem in approximating the full space problem in LancBiO.
@@
for $s,t=1,2,\dots,m$ and
$\varepsilon_{j}^{h}:=\max_{1\leq s,t\leq j}\varepsilon_{st}^{h},$
The following proposition characterizes that the dynamic subspace constructed in Algorithm 2 within an epoch is indeed an approximate Krylov subspace.
@@
$\mathrm{span}(Q_{j+1})\subseteq\mathrm{span}\left\{A_{1}^{a_{1}}A_{2}^{a_{2}}\cdots A_{j}^{a_{j}}\bar{b}\,\left|\!\begin{array}[]{c}a_{s}=0\ \text{\emph{or}}\ 1,\\ s=1,2,\ldots,j\\ \end{array}\!\right.\!\right\}.$
Specifically, when $A_{1}=A_{2}=\cdots=A_{j}=A$ and $Q_{j+1}$ is of full rank,
$\mathrm{span}(Q_{j+1})=\mathcal{K}_{j+1}\left({A,\bar{b}}\right).$
Denote
$A^{*}=\nabla_{yy}^{2}g(x,y^{*})\quad\text{and}\quad b^{*}=\nabla_{y}f(x,y^{*}).$
Notice that the dynamic Lanczos process in Algorithm 2 centers on $A_{j}$ instead of $A^{*}_{j}$. The subsequent lemma interprets the perturbation analysis for the dynamic Lanczos process in terms of $A_{j}$, which satisfies an approximate 3-term recurrence with a perturbation term $\delta Q$.
@@
$A_{j}^{*}Q_{j}=Q_{j}T_{j}+\beta_{j+1}q_{j+1}e_{j}^{\top}+\delta Q_{j}$
$T_{j}=\begin{pmatrix}\alpha_{1}&\beta_{2}&&&&\\ \beta_{2}&\alpha_{2}&\beta_{3}&&&\\ &\beta_{3}&\ddots&\ddots&\\ &&\ddots&\ddots&\beta_{j}&\\ &&&\beta_{j}&\alpha_{j}&\\ \end{pmatrix}$
and $\left\|{\delta q_{j}}\right\|\leq L_{gyy}\varepsilon_{j}$.
@@ 3.2 Convergence Analysis
To guarantee the stable behavior of the dynamic process, three mild assumptions are needed.
@@
The initialization of $y_{1}$ in Algorithm 2 satisfies
$\left\|y_{1}-y_{1}^{*}\right\|\leq\frac{\sqrt{3}\mu_{g}}{8\left(m+1\right)^{3}L_{gyy}}.$
Similar initialization refinement is used in (Hao et al., 2024), which can be achieved by implementing several gradient descent steps for the smooth and strongly convex lower-level problem.
Assumption 3.10.
For each epoch, the following inequality holds,
$\varepsilon_{m}^{h}\leq$ $~{}\lambda\left(1+\frac{L_{gx}}{\mu_{g}}\right)\sum_{j=1}^{m-1}{\left\|\widetilde{\nabla}\varphi\left(x_{mh+j}\right)\right\|}$ (12)
$+\left\|y_{mh+m}-y_{mh+m}^{*}\right\|.$
Assumption 3.11.
There exists a constant $C_{gy}>0$ such that the iterates $(x_{k+1},y_{k})$ generated by Algorithm 2 satisfy $\left\|{\nabla_{y}g(x_{k+1},y_{k})}\right\|\leq C_{gy}$.
3.10 is mild since the first term on the right-hand side of (12) contains all upper-level steps $\{\lambda\widetilde{\nabla}\varphi\left(x_{mh+j}\right)\}_{j=1}^{m-1}$ within one epoch. 3.11 is reasonable in practice since we initialize $y_{1}\approx y^{*}_{1}$ and consider the descent properties of $\left\|{y_{k}-y_{k}^{*}}\right\|$ outlined in Appendix H.
Denote the residual
$\bar{r}_{0}$ $:=\bar{b},$
$\bar{r}_{j}$ $:=\bar{b}-A_{j}^{*}\Delta v_{j},\quad j=1,2,\ldots,m.$
The following lemma reveals that the dynamic process yields an improved solution for the subproblem (10).
@@
Suppose Assumptions 3.1, 3.2, 3.3, 3.6, 3.9, 3.10, and 3.11 hold. Within each epoch, we set the step size $\theta\sim\mathcal{O}({1}/{m^{4}})$ a constant for $y$ and the step size for $x$ as zero in the first $m_{0}\sim\mathcal{O}(1)$ steps, and the others as an appropriate constant $\lambda\sim\mathcal{O}({1}/{m^{4}})$, then we have the following inequality,
@@
Suppose Assumptions 3.1, 3.2, 3.3, 3.6, 3.9, 3.10 and 3.11 hold. Within each epoch, we set the step size as $\theta\sim\mathcal{O}({1}/{m^{4}})$ a constant for $y$ and the step size for $x$ as zero in the first $m_{0}$ steps, and the others as a constant $\lambda\sim\mathcal{O}({1}/{m^{4}})$, then the iterates $\{x_{k}\}$ generated by Algorithm 2 satisfy
$\frac{m}{K\left(m-m_{0}\right)}\sum_{\footnotesize\begin{subarray}{c}k=0,\\ (k\,\emph{{mod}}\,m)>m_{0}\end{subarray}}^{K}{\left\|\nabla\varphi\left(x_{k}\right)\right\|^{2}}\leq\mathcal{O}\left(\frac{m\lambda^{-1}}{K\left(m-m_{0}\right)}\right),$
where $m_{0}\sim\mathcal{O}(\log m)$ is a constant and $m$ is the subspace dimension.
@@
4 Numerical Experiments
In this section, we conduct experiments in the deterministic setting to empirically validate the performance of the proposed algorithms. We test on a synthetic problem and two deep learning tasks. The selection of parameters and more details of the experiments are deferred to Appendix I. We have made the code publicly available on https://github.com/UCAS-YanYang/LancBiO.
Synthetic problem: We concentrate on a synthetic scenario in bilevel optimization (1) with $d_{x}=d_{y}=d$ and
@@ 4 Numerical Experiments
$g(x,y):=$ $~{}c_{2}\sum_{i=1}^{d}{\sin(x_{i}+y_{i})}+\log\left({\sum_{i=1}^{d}{e^{x_{i}y_{i}}}}\right)$
$+\frac{1}{2}y^{\top}\left({D_{3}+G}\right)y.$
@@
It can be seen from Figure 8 that LancBiO achieves the final accuracy the fastest, which benefits from the more accurate $v^{*}$ estimation. Figure 5 illustrates how variations in $m$ and $I$ influence the performance of LancBiO and AmIGO, tested across a range from $10$ to $150$ for $m$, and from $2$ to $10$ for $I$. For clarity, we set the seed of the experiment at $4$, and present typical results to encapsulate the observed trends. It is observed that the increase of $m$ accelerates the decrease in the residual norm, thus achieving better convergence of the hyper-gradient, which aligns with the spirit of the classic Lanczos process. Under the same outer iterations, to attain a comparable convergence property, $I$ for AmIGO-CG should be set to $10$. Furthermore, given that the number of Hessian-vector products averages at $(1+\frac{1}{m})$ per outer iteration for LancBiO, whereas AmIGO requires $I\geq 2$, it follows that LancBiO is more efficient.
Data hyper-cleaning on MNIST: The data hyper-cleaning task (Shaban et al., 2019), conducted on the MNIST dataset (LeCun et al., 1998), aims to train a classifier in a corruption scenario, where the labels of the training data are randomly altered to incorrect classification numbers at a certain probability $p$, referred to as the corruption rate.
The results are presented in Figure 6 and Table 1. Note that LancBiO is crafted for approximating the Hessian inverse vector product $v^{*}$, while the two solid methods, TTSA and stocBiO are not. Consequently, with respect to the residual norm of the linear system, i.e., $\left\|{A_{k}v_{k}-b_{k}}\right\|$, we only compare the results with AmIGO-GD, AmIGO-CG and SOBA. Observe that the proposed subspace-based LancBiO achieves the lowest residual norm and the best test accuracy, and subBiO is comparable to the other algorithms. Specifically, in Figure 6, the efficiency of LancBiO stems from its accurate approximation of the linear system. Additionally, while AmIGO-CG is also adept at approximating $v^{*}$, the green line in Figure 6 indicates that it tends to yield results with higher variance.
@@
Logistic regression on $20$Newsgroup: Consider the hyper-parameters selection task on the $20$Newsgroups dataset (Grazzi et al., 2020), which contains $c=20$ topics with around $18000$ newsgroup posts represented in a feature space of dimension $l=130107$. The goal is to simultaneously train a linear classifier $w$ and determine the optimal regularization parameter $\zeta$. As shown in Figure 11, AmIGO-CG exhibits slightly better performance in reducing the residual norm. Nevertheless, under the same time, LancBiO implements more outer iterations to update $x$, which optimizes the hyper-function more efficiently.
@@
Generally, for standard linear systems, the traditional Lanczos process is recognized for its efficiency and versatility over gradient descent methods. LancBiO, in a sense, reflects this principle within the context of bilevel optimization, underscoring the effectiveness of the dynamic Lanczos-aided approach.
5 Conclusion and Perspective
This paper presents a novel approach to bilevel optimization through the dynamic Lanczos process to approximate Krylov subspaces, to address the computational challenges inherent in calculating the hyper-gradient. The theoretical analysis and empirical validation of LancBiO underscore its reliability and effectiveness. The proposed framework is limited to the deterministic setting. The incorporation of stochastic subspace techniques could unlock better potential in wider applications.

### v2

@@ 2.3 Relation to existing algorithms
$v_{k}$ $=v_{k-1}-\eta\left(A_{k}v_{k-1}-b_{k}\right)=(I-\eta A_{k})v_{k-1}+\eta b_{k},$
while the proposed SubBiO constructs a two-dimensional subspace, $\mathcal{S}_{k}=\operatorname{span}\left\{b_{{k}},(I-\eta A_{{k}})v_{k-1}\right\}$ defined in (7). It is worth noting that the updated $v_{k}$ in SOBA and FSLA belongs to the subspace $\mathcal{S}_{k}$. Furthermore, in the sense of solving the two-dimensional subproblem (6), SubBiO selects the optimal solution $v$ in the subspace.
In addition, if the subspace dimension $m$ is set to one, LancBiO is simplified to a scenario in which one conjugate gradient (CG) step with the warm start mechanism is performed in each outer iteration, which exactly recovers the algorithm AmIGO-CG (Arbel and Mairal, 2022) with one inner iteration to the update of $v$. Alternatively, if the step size $\lambda$ in Algorithm 2 is set to 00 within each $m$-steps (i.e., only inner iterations are invoked), LancBiO reduces to the algorithm AmIGO-CG (Arbel and Mairal, 2022) with $m$ inner iterations.
@@ 3 Theoretical Analysis
The Lipschitz properties of $f,g$ and the strong convexity of the lower-level problem revealed by the above assumptions are standard in bilevel optimization (Ghadimi and Wang, 2018; Chen et al., 2021; Ji et al., 2021; Khanduri et al., 2021; Arbel and Mairal, 2022; Chen et al., 2022; Dagréou et al., 2022; Li et al., 2022; Ji et al., 2022; Hong et al., 2023). These assumptions ensure the smoothness of $\varphi$ and $y^{*}$; see the following results (Ghadimi and Wang, 2018).
@@
Under the Assumptions 3.2 and 3.3, $y^{*}(x)$ is ${L_{gx}}/{\mu_{g}}$Lipschitz continuous, i.e., for any $x_{1},x_{2}\in\mathbb{R}^{d_{x}}$, $\left\|{y^{*}(x_{1})-y^{*}(x_{2})}\right\|\leq\frac{L_{gx}}{\mu_{g}}\left\|{x_{1}-x_{2}}\right\|$.
@@
Under the Assumptions 3.1, 3.2 and 3.3, the hyper-gradient $\nabla\varphi(\cdot)$ is $L_{\varphi}$-Lipschitz continuous, i.e., for any $x_{1},x_{2}\in\mathbb{R}^{d_{x}}$, $\left\|{\nabla\varphi(x_{1})-\nabla\varphi(x_{2})}\right\|\leq L_{\varphi}\left\|{x_{1}-x_{2}}\right\|$, where $L_{\varphi}>0$ is defined in Appendix E.
@@
3.6, commonly adopted in (Ghadimi and Wang, 2018; Ji et al., 2021; Liu et al., 2022; Kwon et al., 2023), is helpful in ensuring the stable behavior of the dynamic Lanczos process; see Section 3.1.
@@ 3.1 Subspace Properties in Dynamic Lanczos Process
In view of the inherent instability of the Lanczos process (Paige, 1980; Meurant and Strakoš, 2006) and the evolution of the Hessian $\{A_{k}\}$ and the gradient $\{b_{k}\}$ in LancBiO, the analysis of the constructed subspaces is intricate. Based on the existing work (Paige, 1976; 1980; Greenbaum, 1997), this subsection sheds light on the analysis of the subspaces and the effectiveness of the subproblem in approximating the full space problem in LancBiO.
@@
for $s,t=1,2,\dots,m$, and $\varepsilon_{j}^{h}:=\max_{1\leq s,t\leq j}\varepsilon_{st}^{h}$, serving as the accumulative difference. For brevity, we omit the superscript where there is no ambiguity, and we are slightly abusing of notation that at the current epoch, $\{A_{mh+j}\}$ and $\{b_{mh+j}\}$ are simplified by $\{A_{j}\}$ and $\{b_{j}\}$ for $j=1,\dots,m$. In addition, the approximations in the residual system (8) are simplified by $\bar{v}$ and $\bar{b}:=b_{1}-A_{1}\bar{v}$.
The following proposition demonstrates that the dynamic subspace constructed in Algorithm 2 within an epoch is indeed an approximate Krylov subspace.
@@
$\mathrm{span}(Q_{j+1})\subseteq\mathrm{span}\left\{A_{1}^{a_{1}}A_{2}^{a_{2}}\cdots A_{j}^{a_{j}}\bar{b}\,\ \big{|}\ a_{s}=0\ \text{\emph{or}}\ 1,\ s=1,2,\ldots,j\right\}.$
Specifically, when $A_{1}=A_{2}=\cdots=A_{j}=A$ and $Q_{j+1}$ is of full rank, $\mathrm{span}(Q_{j+1})=\mathcal{K}_{j+1}\left({A,\bar{b}}\right)$.
Denote $A^{*}=\nabla_{yy}^{2}g(x,y^{*})$ and $b^{*}=\nabla_{y}f(x,y^{*})$. Notice that the dynamic Lanczos process in Algorithm 2 centers on $A_{j}$ instead of $A^{*}_{j}$. The subsequent lemma interprets the perturbation analysis for the dynamic Lanczos process in terms of $A^{*}_{j}$, which satisfies an approximate three-term recurrence with a perturbation term $\delta Q$.
@@
Suppose Assumptions 3.1 to 3.3 hold. The dynamic Lanczos process in Algorithm 2 with normalized $q_{1}$ satisfies
$A_{j}^{*}Q_{j}=Q_{j}T_{j}+\beta_{j+1}q_{j+1}e_{j}^{\top}+\delta Q_{j},\ \ \text{for}\ j=1,2,\dots,m,$
where $Q_{j}=\left[q_{1},q_{2},\dots,q_{j}\right]$, $\delta Q_{j}=\left[\delta q_{1},\delta q_{2},\dots,\delta q_{j}\right]$ with $\left\|{\delta q_{j}}\right\|\leq L_{gyy}\varepsilon_{j}$, and $T_{j}$ is a $j\times j$ symmetric tridiagonal matrix with diagonal elements $\left\{{\alpha_{1},\ldots,\alpha_{j}}\right\}$ and subdiagonal elements $\left\{{\beta_{2},\ldots,\beta_{j}}\right\}$.
@@ 3.2 Convergence Analysis
To guarantee the stable behavior of the dynamic process, we need the subsequent assumption.
@@
[…]
Similar initialization refinement is used in (Hao et al., 2024), which can be achieved by implementing several gradient descent steps for the smooth and strongly convex lower-level problem. The following lemma reveals that the dynamic process yields an improved solution for the subproblem (9).
@@
Suppose Assumptions 3.1, 3.2, 3.3, 3.6 and 3.9 hold. Within each epoch, if we set the step size $\theta\sim\mathcal{O}({1}/{m})$ a constant for $y$, and the step size for $x$ as zero in the first $m_{0}\sim\Omega(1)$ steps and the others as an appropriate constant $\lambda\sim\mathcal{O}({1}/{m^{4}})$, we have the following inequality,
@@
Suppose Assumptions 3.1, 3.2, 3.3, 3.6 and 3.9 hold. Within each epoch, if we set the step size $\theta\sim\mathcal{O}({1}/{m})$ a constant for $y$, and the step size for $x$ as zero in the first $m_{0}$ steps and the others as an appropriate constant $\lambda\sim\mathcal{O}({1}/{m^{4}})$, the iterates $\{x_{k}\}$ generated by Algorithm 2 satisfy
$\frac{m}{K\left(m-m_{0}\right)}\sum_{\begin{subarray}{c}k=0,\\ (k\,\emph{{mod}}\,m)>m_{0}\end{subarray}}^{K}{\left\|\nabla\varphi\left(x_{k}\right)\right\|^{2}}=\mathcal{O}\left(\frac{m\lambda^{-1}}{K\left(m-m_{0}\right)}\right),$
where $m_{0}\sim\Omega(\log m)$ is a constant and $m$ is the subspace dimension.
@@
4 Numerical Experiments
In this section, we conduct experiments in the deterministic setting to empirically validate the performance of the proposed algorithms. We test on a synthetic problem and two deep learning tasks. The selection of parameters and more details of the experiments are deferred to Appendix I. We have made the code available on https://github.com/UCAS-YanYang/LancBiO.
@@ 4 Numerical Experiments
F2SA $89.78\pm 0.1969$ $0.40\pm 0.0073$ $88.65\pm 0.2828$ $0.51\pm 0.0055$
HJFBiO $90.21\pm 0.2027$ $0.37\pm 0.0048$ $89.30\pm 0.3594$ $0.43\pm 0.0040$
Data hyper-cleaning on three datasets: The data hyper-cleaning task (Shaban et al., 2019) aims to train a classifier in a corruption scenario, where the labels of the training data are randomly altered to incorrect classification numbers at a probability $p$, referred to as the corruption rate. The results on the MNIST dataset are presented in Figure 4 and Table 1. Note that LancBiO is crafted for approximating the Hessian inverse vector product $v^{*}$, while the two solid methods, TTSA and stocBiO are not. Consequently, with respect to the residual norm of the linear system, i.e., $\left\|{A_{k}v_{k}-b_{k}}\right\|$, we only compare the results with AmIGO-GD, AmIGO-CG, and SOBA. Observe that the proposed subspace-based LancBiO achieves the lowest residual norm and the best test accuracy, and subBiO is comparable to the other algorithms. Specifically, in Figure 4, the efficiency of LancBiO stems from its accurate approximation of the linear system. Additionally, while AmIGO-CG is also adept at approximating $v^{*}$, the results in Table 1 indicate that it tends to yield higher variance. Moreover, algorithms are also evaluated on the Fashion-MNIST and Kuzushiji-MNIST datasets; see Figure 10 and Figure 11, respectively. The proposed LancBiO performs better than other algorithms and showcases robustness across various datasets.
Synthetic problem: We concentrate on a synthetic bilevel optimization (1) with $d_{x}=d_{y}=d$ and
@@
$g(x,y):=$ $~{}c_{2}\sum_{i=1}^{d}{\sin(x_{i}+y_{i})}+\log\left({\sum_{i=1}^{d}{e^{x_{i}y_{i}}}}\right)+\frac{1}{2}y^{\top}\left({D_{3}+G}\right)y.$
@@
It can be seen from Figure 12 that LancBiO achieves the final accuracy the fastest, which benefits from the more accurate $v^{*}$ estimation. Figure 5 illustrates how variations in $m$ and $I$ influence the performance of LancBiO and AmIGO, tested across a range from $10$ to $150$ for $m$, and from $2$ to $10$ for $I$. For clarity, we set the seed of the experiment at $4$, and present typical results to encapsulate the observed trends. It is observed that the increase of $m$ accelerates the decrease in the residual norm, thus achieving better convergence of the hyper-gradient, which aligns with the spirit of the classic Lanczos process. Under the same outer iterations, to attain a comparable convergence property, $I$ for AmIGO-CG should be set to $10$. Furthermore, given that the number of Hessian-vector products averages at $(1+1/m)$ per outer iteration for LancBiO, whereas AmIGO involves $I\geq 2$ calculations, it follows that LancBiO is more efficient. Moreover, to illustrate how the methods scale with increasing dimensions, we present the convergence time and the final upper-level value under different problem dimensions $d=10^{i},i=1,2,3,4$ in Table 2. The results demonstrate the proposed methods maintain decent performance across different problem dimensions.
Figure 6: Comparison of the bilevel algorithms on the hyper-parameters selection task. Left: validation accuracy; Center: validation loss; Right: residual norm of the linear system, $\left\|{A_{k}v_{k}-b_{k}}\right\|$.
@@
Logistic regression on $20$Newsgroup: Consider the hyper-parameter selection task on the $20$Newsgroups dataset (Grazzi et al., 2020). The goal is to train a linear classifier $w$ and determine the optimal regularization parameter $\zeta$. As shown in Figure 6, AmIGO-CG exhibits slightly better performance in reducing the residual norm. Nevertheless, under the same time, LancBiO implements more outer iterations to update $x$, which optimizes the hyper-function more efficiently.
@@
Generally, to solve standard linear systems, the Lanczos process is recognized for its efficiency and versatility over gradient descent methods. LancBiO, in a sense, reflects this principle in the context of bilevel optimization, underscoring the effectiveness of the dynamic Lanczos-aided approach.
Acknowledgments
Bin Gao was supported by the Young Elite Scientist Sponsorship Program by CAST. Ya-xiang Yuan was supported by the National Natural Science Foundation of China (grant No. 12288201). The authors are grateful to the Program and Area Chairs and Reviewers for their valuable comments and suggestions.
Appendix